OpenStudy (anonymous):

A tank is holding 60 liters of a chemical solution. Currently, the solution has a strength of 30%. How much 70% solution must be added to the tank to get a final strength of 40%?

OpenStudy (anonymous):

man those hits just keep coming

OpenStudy (anonymous):

haha I'm sorry but I'm serious when I say I am horrible at math ..

OpenStudy (anonymous):

you have 60 liters that is 30% solution, so at the the start you have \(.3\times 6=1.8\) liters of solution

OpenStudy (anonymous):

if you add say \(x\) liters of 70% solution, it will have \(.7x\) liters of solution, and you will have \(60+x\) liters total, that you want to be 40% solutions

OpenStudy (anonymous):

therefore you can solve \[1.8+.7x=.4(60+x) \] for \(x\)

OpenStudy (anonymous):

easier to multiply by 10 to get rid of the annoying decimal, and solve instead \[18+7x=4(60+x)\]

OpenStudy (anonymous):

still need help, or are you good from here?

OpenStudy (anonymous):

74?

OpenStudy (anonymous):

yup guess you got this

OpenStudy (anonymous):

haha